OpenStudy (anonymous):

whats the simplified form of the square root of 1044

OpenStudy (anonymous):

\[9\sqrt{29}\]

OpenStudy (anonymous):

\[\sqrt{36 * 29} = 6\sqrt{29}\]

OpenStudy (anonymous):

Yeah, what he said.

OpenStudy (anonymous):

thank youu

OpenStudy (anonymous):

do like that(1044)^1/2.and hit the calculator

OpenStudy (anonymous):

lol unless you have a ti 89, it will give you a decimal, i think she asking for the exact number

OpenStudy (anonymous):

you can round off

OpenStudy (anonymous):

and if the teacher asks for a exact answer?

OpenStudy (anonymous):

\[\sqrt{1044} = \sqrt{36*29} = \sqrt{6^2*29} = 6\sqrt{29}\]